# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

### v4

@@ Abstract
The effectiveness of deep learning models in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper addresses this problem by leveraging bias amplification with generated synthetic data only: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing, exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which fully replace the original training set for learning an effective bias amplifier model to be subsequently incorporated into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples’ memorization in learning auxiliary models, typical of this type of technique, our proposed method outperforms the current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
@@ 2 Related Work
The closest work employing generative models for model debiasing is Jung et al. jung2023fighting , where a debiasing method exploiting a GAN 10.1145/3422622 is introduced, which is trained for style-transfer between several biased appearances with different objectives. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations. Similar strategies, where bias-conflicting samples are augmented to support model debiasing, can be found in a few other works biasadv ; hwang2022selecmix ; dataaug1 . In dataaug1 , the problem of bias-conflicting memorization is stressed, proposing a bias-unsupervised strategy to augment and increase the diversity of bias-conflicting features, exploiting a bias amplifier analogously to nam2020learning for their identification, while improving the generalization of a target debiased model. In biasadv , an intentionally biased auxiliary model is exploited, which is trained on the biased dataset, supplementing bias-conflicting samples for the target debiased model. Here, images are generated by solving an optimization problem consisting of attacking the auxiliary model predictions while preserving the ones of the target debiased model. In hwang2022selecmix , bias-conflicting samples are again augmented exploiting the popular mix-up approach zhang2018mixupempiricalriskminimization . Specifically, an auxiliary model is employed to amplify easy-to-learn features, and mix-up pairs with either the same ground truth label, but different biased features (evaluated with clustering in the feature space of an intentionally biased model), or samples with different biased features (i.e., belonging to different clusters), but showing the same ground truth label.
@@ 3.1 Problem Setting
Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution. Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e., possible biases such as context, appearance, acquisition noise, viewpoint, etc.. In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not present the same biases. In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting kim2024training . If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization. This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity kim2021biaswap ; Sagawa*2020Distributionally ; nahon2023mining ), which is dominating over the true target distribution semantics.
It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation d2024openbias . For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in d2024openbias ; kim2024training . While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature nam2020learning ), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
@@ 3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an auxiliary intentionally-biased model, named here as Bias Amplifier (BA). This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
@@ 3.3.1 Recipe I: 2-step debiasing
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 2(a).
@@ 4 Experiments
In this Section, we report the description of the benchmark datasets employed, followed by a discussion of the obtained results and a thorough ablation analysis.
In the Appendix, we report the general implementation (Sec. A.2) and training details, including the implementation details of Recipe I (Sec. A.3) and Recipe II (Sec. A.4), and more examples of the generated images (Sec. A.6).
@@ 4.2.2 Classification accuracy
Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in Sagawa*2020Distributionally ; Lee_Park_Kim_Lee_Choi_Choo_2023 . For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following kim2021biaswap ; li2022discover . BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It is worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
Method Unsup Waterbirds BAR BFFHQ ImageNet-A
WGA Avg. Avg. Confl. Avg.
[…]
@@
[…]
@@
[…]
CDvG+LfF jung2023fighting $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
[…]
Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1). Notably, in Waterbirds, DDB’s WGA is higher than all the other state-of-the-art unsupervised methods’ accuracies, as well as of that of the supervised G-DRO Sagawa*2020Distributionally (91.56% vs. 91.40% of G-DRO). DDB-I is providing slightly worse performance than DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN li2022discover ), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to the state-of-the-art ETF-Debias wang2024navigate (73.60%).
@@
[…]
Table 2: Comparison of deep learning methods on the UrbanCars biased dataset. ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better).
@@ 4.3 Ablation analysis
$w$ 0 1 2 3 5
WGA 87.85 90.81 89.25 89.56 88.32
Table 3: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds, employing Recipe I.
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Table 1, the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like in BFFHQ ($\rho=0.995$).
@@ 5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased model. The usage of synthetic bias-aligned data prevents the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. DDB results effective and versatile, acting as a plug-in for unsupervised debiasing methods. In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes. leading to accuracies outperforming state-of-the-art works by a margin on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say . Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable to be adopted as an effective general-purpose standard training procedure.
Limitations. DDB’s main limitations rely on the well-known high computational complexity of diffusion models. As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM. Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM. Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.

### v5

@@ Abstract
The effectiveness of deep learning models in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper addresses this problem by leveraging bias amplification with generated synthetic data only: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing, exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which fully replace the original training set for learning an effective bias amplifier model to be subsequently incorporated into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples’ memorization in learning auxiliary models, typical of this type of technique, our proposed method outperforms the current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models. Code is available at https://github.com/Malga-Vision/DiffusingDeBias.
@@ 2 Related Work
The closest work employing generative models for model debiasing is Jung et al. jung2023fighting , where a debiasing method exploiting a GAN 10.1145/3422622 is introduced, which is trained for style-transfer between several biased appearances with different objectives. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations. Similar strategies, where bias-conflicting samples are augmented to support model debiasing, can be found in a few other works biasadv ; hwang2022selecmix ; dataaug1 . In dataaug1 , the problem of bias-conflicting memorization is stressed, proposing a bias-unsupervised strategy to augment and increase the diversity of bias-conflicting features, exploiting a bias amplifier analogously to nam2020learning for their identification, while improving the generalization of a target debiased model. In biasadv , an intentionally biased auxiliary model is exploited, which is trained on the biased dataset, supplementing bias-conflicting samples for the target debiased model. Here, images are generated by solving an optimization problem consisting of attacking the auxiliary model predictions while preserving the ones of the target debiased model. In hwang2022selecmix , bias-conflicting samples are again augmented exploiting the popular mix-up approach zhang2018mixupempiricalriskminimization . Specifically, an auxiliary model is employed to amplify easy-to-learn features, and mix-up pairs with either the same ground truth label, but different biased features (evaluated with clustering in the feature space of an intentionally biased model), or samples with different biased features (i.e., belonging to different clusters), but showing the same ground truth label. Other works have also explored generative models for debiasing. For instance, BiaSwap kim2021biaswap utilizes image translation techniques to swap bias attributes between images, aiming to create more balanced training data. However, methods that attempt to create synthetic bias-conflicting samples kim2021biaswap ; biasadv may require high semantic fidelity, while our method creates a functionally biased dataset, which we believe to be a simpler and more robust objective.
@@ 3.1 Problem Setting
[…]
@@ 3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an auxiliary intentionally-biased model, named here as Bias Amplifier (BA). This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively. This section describes the two debiasing recipes we employ in this work to demonstrate the effectiveness of our BA as a plug-in for unsupervised model debiasing.
@@ 3.3.1 Recipe I: 2-step debiasing
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma$$\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution ($\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})$} in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 2(a).
@@ 4 Experiments
In this Section, we report the description of the benchmark datasets employed, followed by a discussion of the obtained results and a thorough ablation analysis. In the Appendix, we report the general implementation (Sec. A.2) and training details, including the implementation details of Recipe I (Sec. A.3) and Recipe II (Sec. A.4), more examples and analysis on the generated images (Sec. A.6), and additional debiasing results on less biased settings for Waterbirds plus an alternative version of BAR having known $\rho$ and bias-annotations.
@@ 4.2.2 Classification accuracy
Tables 1(a), 1(b), 2(a), 2(b), and 3 report the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered benchmark datasets. We dedicate a separate table for each dataset, as not all methods share the same benchmarks. We report Worst Group Accuracy (WGA) for Waterbirds, as in Sagawa*2020Distributionally ; Lee_Park_Kim_Lee_Choi_Choo_2023 . For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following kim2021biaswap ; li2022discover . BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It is worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
Table 1: Results of DDB Recipes I and II on Waterbirds (left) and BAR (right). Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
(a) Debiasing results on Waterbirds
Method Unsup Waterbirds
WGA
ERM $62.60\text{\scriptsize{$\pm$ 0.30}}$
@@
G-DRO Sagawa*2020Distributionally $91.40\text{\scriptsize{$\pm$ 1.10}}$
George sohoni2020no $76.20\text{\scriptsize{$\pm$ 2.00}}$
JTT pmlr-v139-liu21f_JTT $83.80\text{\scriptsize{$\pm$ 1.20}}$
CNC pmlr-v162-zhang22z $88.50\text{\scriptsize{$\pm$ 0.30}}$
B2T+G-DRO kim2024discovering $90.70\text{\scriptsize{$\pm$ 0.30}}$
LfF nam2020learning $78.00$
CDvG+LfF jung2023fighting $84.80$
MoDAD pastore2024lookingmodeldebiasinglens $89.43\text{\scriptsize{$\pm$ 1.69}}$
DDB-II (ours) $\bf 91.56\text{\scriptsize{$\pm$ 0.15}}$
DDB-I (ours) $90.81\text{\scriptsize{$\pm$ 0.68}}$
DDB-I (w/ err. set) $90.34\text{\scriptsize{$\pm$ 0.41}}$
(b) Debiasing results on BAR
Method Unsup BAR
Avg.
ERM $51.85\text{\scriptsize{$\pm$ 5.92}}$
BiaSwap kim2021biaswap $52.40\text{\scriptsize{$\pm$ -}}$
JTT pmlr-v139-liu21f_JTT $68.53\text{\scriptsize{$\pm$ 3.29}}$
LfF nam2020learning $62.98\text{\scriptsize{$\pm$ 2.76}}$
LWBC NEURIPS2022_75004615_LWBC $62.03\text{\scriptsize{$\pm$ 2.76}}$
DebiAN li2022discover $69.88\text{\scriptsize{$\pm$ 2.92}}$
MoDAD pastore2024lookingmodeldebiasinglens $69.83\text{\scriptsize{$\pm$ 0.72}}$
DDB-II (ours) $\bf 72.81\text{\scriptsize{$\pm$ 1.02}}$
DDB-I (ours) $70.40\text{\scriptsize{$\pm$ 1.41}}$
DDB-I (w/ err. set) $70.59\text{\scriptsize{$\pm$ 0.19}}$
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results. Notably, in Waterbirds, DDB’s WGA is higher than all the other state-of-the-art unsupervised methods’ accuracies, as well as of that of the supervised G-DRO Sagawa*2020Distributionally (91.56% vs. 91.40% of G-DRO). DDB-I is providing slightly worse performance than DDB-II, but still better than the state of the art.
Table 2: Results of DDB Recipes I and II on BFFHQ (left) and ImageNet-9/A (right) from kim2021biaswap . Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
(a) Debiasing results on BFFHQ
Method Unsup BFFHQ
Avg. Confl.
ERM $60.13\text{\scriptsize{$\pm$ 0.46}}$
BiaSwap kim2021biaswap $58.87\text{\scriptsize{$\pm$ -}}$
JTT pmlr-v139-liu21f_JTT $62.20\text{\scriptsize{$\pm$ 1.34}}$
LfF nam2020learning $62.97\text{\scriptsize{$\pm$ 3.22}}$
ETF-Debias wang2024navigate $73.60\text{\scriptsize{$\pm$ 1.22}}$
@@
CDvG+LfF jung2023fighting $62.20\text{\scriptsize{$\pm$0.45}}$
DebiAN li2022discover $62.80\text{\scriptsize{$\pm$ 0.60}}$
MoDAD pastore2024lookingmodeldebiasinglens $68.33\text{\scriptsize{$\pm$ 2.89}}$
DDB-II (ours) $\bf 83.15\text{\scriptsize{$\pm$ 1.76}}$ $70.93\text{\scriptsize{$\pm$ 0.14}}$
DDB-I (ours) $81.27\text{\scriptsize{$\pm$ 0.88}}$ $\bf 74.67\text{\scriptsize{$\pm$ 2.37}}$
DDB-I (w/ err. set) $82.44\text{\scriptsize{$\pm$ 0.64}}$ $71.40\text{\scriptsize{$\pm$ 0.92}}$
(b) Debiasing results on ImageNet-9/A
Method Unsup ImageNet-A
Avg.
ERM $30.30$
LWBC NEURIPS2022_75004615_LWBC $35.97\text{\scriptsize{$\pm$0.49}}$
CDvG+LfF jung2023fighting $34.60$
DDB-II (ours) $37.53\text{\scriptsize{$\pm$0.82}}$
DDB-I (ours) $\bf 39.80\text{\scriptsize{$\pm$0.50}}$
DDB-I (w/ err. set) $38.12\text{\scriptsize{$\pm$0.96}}$
In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN li2022discover ), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to the state-of-the-art ETF-Debias wang2024navigate (73.60%).
Table 3: Comparison of deep learning methods on the UrbanCars biased dataset. ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better). GroupDRO and LLE are separated as they rely on bias information.
@@
DDB-I (ours) $86.39\text{\scriptsize{$\pm$ 0.74}}$ $\bf 1.85\text{\scriptsize{$\pm$ 3.21}}$ $\bf 0.52\text{\scriptsize{$\pm$ 1.38}}$ $\bf 0.12\text{\scriptsize{$\pm$ 1.56}}$
DDB-II (ours) $\bf 98.56\text{\scriptsize{$\pm$ 0.92}}$ $2.30\text{\scriptsize{$\pm$ 0.60}}$ $11.10\text{\scriptsize{$\pm$ 1.20}}$ $46.70\text{\scriptsize{$\pm$ 2.42}}$
@@ 4.3 Ablation analysis
Table 4: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds. For both experiments, we employ Recipe I.
@@
$w$ 0 $\mathbf{1}$ 22 33 55
WGA $87.85$ $\mathbf{90.81}$ $89.25$ $89.56$ $88.32$
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Tables 1(a), 1(b), 2(a), and 2(b), the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications is always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like in BFFHQ ($\rho=0.995$).
@@
Effectiveness of Synthetic Data for BA Training. A core claim of our work is that training the Bias Amplifier on synthetic data prevents the memorization of real bias-conflicting samples. To validate this, we compare our standard BA (trained on synthetic data) with a BA trained on the original real training data within our two recipes. Using the real dataset significantly degrades performance. For Recipe I, the BA trained on real data for 50 epochs memorized all conflicting samples, leading to empty pseudo-label groups. Training for just one epoch (as in JTT pmlr-v139-liu21f_JTT ) yielded a WGA of only 79.43%. For Recipe II, using a frozen BA trained on real data resulted in a WGA of 78.45%, far below our proposed method and in line with the original LfF performance. These results empirically confirm that using synthetic data is crucial for creating an effective BA.
@@ 5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased model. The usage of synthetic bias-aligned data is empirically shown to prevent the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. Our experiments confirm that a bias amplifier trained on real data fails at this task, while our approach creates a robust and effective supervisory signal. DDB is versatile, acting as a plug-in for unsupervised debiasing methods, since its Bias Amplifier can be incorporated into several debiasing approaches. In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes, outperforming state-of-the-art works on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say . Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable to be adopted as an effective general-purpose standard training procedure.
Limitations. DDB’s main limitations stem from the high computational cost of training diffusion models. As a trade-off between efficiency and generation quality, we reduce the training image resolution to $64\times 64$; nonetheless, training our CDPM still requires $\sim$14 hours on an NVIDIA A30 GPU with 24 GB of VRAM for the largest considered dataset (BFFHQ, 19,200 images). Our ablations show, however, that this cost can be substantially reduced by training for fewer iterations or at lower resolutions (e.g., 3.2 hours for Waterbirds at $32\times 32$) with only a minor impact on final performance. Importantly, this computational cost should be weighed against the significant, often impractical, burden of manually annotating a bias-aware validation set, which many alternative methods require [34, 43]. Another limitation inherited from diffusion models is that generation quality may decrease if the training dataset is too small, potentially limiting DDB’s applicability in very small-scale settings. Finally, while we have benchmarked DDB on a wide range of standard datasets, the bias mitigation community currently lacks million-scale datasets with appropriate bias annotations for large-scale evaluation, which remains an important direction for future research.
Acknowledgements.
We acknowledge the financial support from PNRR MUR Project PE0000013 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU, CUP J33C24000430007, CUP C63C22000770006.